OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9.

OpenStudy (anonymous):

i'll let @Fgcbear16 do the honors.

OpenStudy (anonymous):

It's a repeat.

OpenStudy (anonymous):

y intercept is given: (0,9) and slope = -6 So, y - 9 = -6(x - 0) y = -6x + 9

OpenStudy (anonymous):

since it is slope intercept y=mx+b m=-6 and b=9 so y=-6x+9